OpenStudy (anonymous):

2. The electrolyte in automobile lead storage batteries is a 3.75 M sulfuric acid solution that has a density of 1.230 g/mL. Calculate the mass percent, molality, and mol fraction of the sulfuric acid. i am so lost dont know how to get the kg of solvent

OpenStudy (abmon98):

\[1.230 g/ml*1000ml/L=1.230*10^3 g/L\] \[3.75 mol*((2*1)+(32*1)+(16*4)) g H2SO4/1 mol=368 g H2SO4\] \[1230.0 g Solution-368 g (H2SO4)=862 g H2O \] \[Mass Percentage of H2SO4=(mass of H2SO4/mass of solution)*100\] \[Mass Percentage of H2SO4=(368.0 g /1230.0g)*100%=29.9% \]

OpenStudy (abmon98):

\[Molality=moles of solute/Kilograms of Solvent \]\[Molality=3.75 mol H2SO4/862*1kg H2O/1000g H2O \]